a triangle has a hypotenuse of 25 inches, its height is 12 inches, what is the perimeter of the triangle?
Answer by stanbon(75887)

a triangle has a hypotenuse of 25 inches, its height is 12 inches, what is the perimeter of the triangle?
-------------------------------
Let the width be "w".
Equation:
25^2 = 12^2 + w^2
625 = 144 + w^2
w^2 = 481
w = 21.93 inches
-----------------------
Perimeter = 2(length + width)
P = 2(25 + 21.93)
P = 93.86 inches
=============================
Cheers,
Stan H.